Of the people in attendance at a recent baseball game
. 1/3 had grandstand tickets,
.1/4 had bleacher tickets, and
. the remaining 11,250 people in attendance had other tickets.
what was the total number of people in attendance at the game?

Answer by checkley79(3341)

1/3X+1/4X+11,250=X
(4+3)/(3*4)X+11,250=X
7/12X+11,250=X
7/12X-X=-11,250
-5X/12=-11,250
=5X=-11,250*12
-5X=-135,700
X=-137,500/-5
X=27,000 ANS. FOR THE TOTAL NUMBER IN ATTENDANCE.
PROOF:
27,000/3+27,000/4+11,250=27,000
9,000+6,750+11,250=27,000
27,000=27,000